OpenStudy (rock_mit182):

a) Show that the eigenvalues of the 2 X 2 matrix A = [a,b];[c,d] are the solutions of the quadratic equation A^2 - tr(A)lambda + det A = 0, where tr(A) is the trace of A.

OpenStudy (rock_mit182):

@amistre64

OpenStudy (amistre64):

i thought AA was just another matrix ... the trace and det are scalars if memory serves also, refresh my memory on the definition of the trace

OpenStudy (amistre64):

sum of the main diag ...

OpenStudy (amistre64):

but then A^2 has me befuddled

OpenStudy (michele_laino):

hint: we can write the subsequent quadratic equation: \[\det \left( {\begin{array}{*{20}{c}} {a - \lambda }&b \\ c&{d - \lambda } \end{array}} \right) = {\lambda ^2} - \left( {a + d} \right)\lambda + ad - bc\]